If

A=[ 1 -3 2
4 1 -1
-3 2 5 ]
and

B=[ 2 1 5
-1 -2 -2
3 1 2 ]

then find (AB)^-1
1
Expert's answer
2020-12-21T18:52:56-0500

Solution:


1) We should multiply both matrix to each other in order to make C matrix:


"AB=A\cdot B=\begin{bmatrix}
 1 & -3 &2\\
 4 & 1&-1\\-3&2&5
\end{bmatrix} \;\cdot\begin{bmatrix}
 2 & 1 &5\\
 -1 & -2&-2\\3&1&2
\end{bmatrix} =\begin{bmatrix}
 11 & 9 &15\\
 4 & 1&16\\7&-2&-9
\end{bmatrix}"


The components of the matrix AB is calculated as:

c11 = a11 · b11 + a12 · b21 + a13 · b31 = 1 · 2 + (-3) · (-1) + 2 · 3 = 2 + 3 + 6 = 11


c12 = a11 · b12 + a12 · b22 + a13 · b32 = 1 · 1 + (-3) · (-2) + 2 · 1 = 1 + 6 + 2 = 9


c13 = a11 · b13 + a12 · b23 + a13 · b33 = 1 · 5 + (-3) · (-2) + 2 · 2 = 5 + 6 + 4 = 15


c21 = a21 · b11 + a22 · b21 + a23 · b31 = 4 · 2 + 1 · (-1) + (-1) · 3 = 8 - 1 - 3 = 4


c22 = a21 · b12 + a22 · b22 + a23 · b32 = 4 · 1 + 1 · (-2) + (-1) · 1 = 4 - 2 - 1 = 1


c23 = a21 · b13 + a22 · b23 + a23 · b33 = 4 · 5 + 1 · (-2) + (-1) · 2 = 20 - 2 - 2 = 16


c31 = a31 · b11 + a32 · b21 + a33 · b31 = (-3) · 2 + 2 · (-1) + 5 · 3 = (-6) - 2 + 15 = 7


c32 = a31 · b12 + a32 · b22 + a33 · b32 = (-3) · 1 + 2 · (-2) + 5 · 1 = (-3) - 4 + 5 = -2


c33 = a31 · b13 + a32 · b23 + a33 · b33 = (-3) · 5 + 2 · (-2) + 5 · 2 = (-15) - 4 + 10 = -9


2) In order to find the inverse of AB, we should use the cofactor method:


I) Find determinant of AB matrix:

Use the triangle's rule to calculate the determinant of the matrix with size 3×3:


det AB = 1360


The determinant of АB is not zero, therefore the inverse matrix (AB)-1 exists. To calculate the inverse matrix find additional minors and cofactors of matrix АB:


  • Find the minor M11 and the cofactor C11. In the matrix, AB crosses out row 1 and column 1.
"M_{11}=\begin{vmatrix}
 1 & 16 \\
 -2 & -9
\end{vmatrix}=23"

"C_{11} = (-1)^{1+1}M_{11} = 23"


  • Find the minor M12 and the cofactor C12. In the matrix, AB crosses out row 1 and column 2.


"M_{12}=\begin{vmatrix}
 4 & 16 \\
 7 & -9
\end{vmatrix}=-148"

  • Find the minor M13 and the cofactor C13. In matrix AB cross out row 1 and column 3.



"M_{13}=\begin{vmatrix}
 4 & 1 \\
 7 & 2
\end{vmatrix}=-15"



  • Find the minor M21 and the cofactor C21. In the matrix, AB crosses out row 2 and column 1.


"M_{21}=\begin{vmatrix}
 9 & 15 \\
 -2 & -9
\end{vmatrix}=-51"


  • Find the minor M22 and the cofactor C22. In a matrix, AB crosses out row 2 and column 2.
"M_{22}=\begin{vmatrix}
 11 & 15 \\
 7 & -9
\end{vmatrix}=-204"

  • Find the minor M23 and the cofactor C23. In matrix AB cross out row 2 and column 3.
"M_{23}=\begin{vmatrix}
 11 & 9 \\
 7 & -2
\end{vmatrix}=-85"

  • Find the minor M31 and the cofactor C31. In matrix AB cross out row 3 and column 1.
"M_{31}=\begin{vmatrix}
 9 & 15 \\
 1 & 16
\end{vmatrix}=129"

  • Find the minor M32 and the cofactor C32. In matrix AB cross out row 3 and column 2.
"M_{32}=\begin{vmatrix}
 11 & 15 \\
 4 & 16
\end{vmatrix}=116"



  • Find the minor M33 and the cofactor C33. In matrix AB cross out row 3 and column 3.
"M_{33}=\begin{vmatrix}
 11 & 9 \\
 4 & 1
\end{vmatrix}=25"

3) Write the matrix of cofactors:


"AB=\begin{bmatrix}
 23 & -148 &-15\\
 51 & 204&85\\129&-116&-25
\end{bmatrix}"


4) Transposed matrix of cofactors:


"AB^T=\begin{bmatrix}
 23 & 51 &129\\
 148 & -204&-116\\-15&85&-25
\end{bmatrix}"


5) Find the inverse matrix:


"AB^{-1}=\frac{AB^T}{det\;(AB)}=\begin{bmatrix}
 \frac{23}{1360} & \frac{3}{80} &\frac{129}{1360}\\
 \frac{37}{340} & -\frac{3}{20}&-\frac{29}{340}\\-\frac{3}{272}&\frac{1}{16}&-\frac{5}{272}
\end{bmatrix}"